OpenStudy (anonymous):

ln(x+5)=ln(x-1)-ln(x+1) solve the logarithmic function algebraically. round the result to three decimal places.

myininaya (myininaya):

hint ln(a)-ln(b)=ln(a/b)

myininaya (myininaya):

then after you use that on the right hand side you just need to set the insides equal

OpenStudy (anonymous):

\[\ln(x+5)=\ln(\frac{x-1}{x+1})\] and therefore the inputs are equal

OpenStudy (anonymous):

what myininay said

myininaya (myininaya):

what satellite said

OpenStudy (anonymous):

lol 500 to go!